# Simple Algorithms for Fully Dynamic Edge Connectivity

### v1

@@ 1.1 Contributions
Our second algorithm improves on the result of [GHN+23] for large edge connectivity, namely when $\lambda_{G}=\omega(\sqrt{n})$. Our result is actually stronger, since it is parametrized by the minimum degree of the graph at the time of the update, $\delta_{G}$, which is only larger than the edge connectivity $\lambda_{G}$ and hence it implies an update time of $\tilde{O}(n^{2}/\lambda^{2}_{G})$. Therefore, Theorem 1.2 improves on [GHN+23] for large edge connectivity, i.e. $\lambda_{G}\geq\omega(\sqrt{n})$, and is in fact the first algorithm to achieve worst-case update time $o(n)$ for large edge connectivity. Moreover, the improvement applies also to dense graphs with small edge connectivity. The improvement is even larger when separating the update and query times, i.e. if the algorithm only needs to report the edge connectivity occasionally.
@@ 1.2 Future Work
The main open question remaining is to achieve worst-case update time $o(n)$ for all edge connectivity values. Prior work achieves this for edge connectivity $\lambda_{G}=o(n^{1/11})$ [Tho07, JST24, dVC25]. Our results complement these results and achieve it for large edge connectivity, namely $\lambda_{G}=\omega(\sqrt{n})$. Perhaps surprisingly, the remaining case is that of intermediate connectivity, $\lambda_{G}\in[n^{1/11},n^{1/2}]$.
@@ 2 Proof of Main Theorems
Our algorithms follow the general recipe of [GHN+23], with the main difference being that our contraction algorithm maintains an explicit KT-style contraction as a weighted graph. [^2] To recover the edges of a minimum cut we can maintain a list of the edges of $G$ that are mapped to each edge of $G^{\prime}$ and report the edges corresponding to the minimum cut in $G^{\prime}$. This both simplifies the algorithm and improves the update time whenever $\delta_{G}\geq\omega(\sqrt{n})$. Specifically, we leverage the $\tau$-star contraction procedure of [AEG+22, KK25], which was originally designed for finding a global minimum cut in the cut-query model. The procedure works as follows. Sample a set of center vertices $R\subseteq V$ by including each vertex $v\in V$ independently with probability $p=O(\log n/\tau)$. Let $H={\{v\in V\setminus R\mid d_{G}(v)\geq\tau\}}$ be the set of non-center vertices with degree at least $\tau$, where throughout $d_{G}(v)$ is the degree of $v$ in $G$. Then, for every $v\in H$, uniformly sample a vertex $r\in N_{G}(v)\cap R$, where $N_{G}(v)$ is the neighborhood of $v$ in $G$, and contract the corresponding edge $(v,r)$, keeping parallel edges, which yields a contracted multigraph $G^{\prime}$. The main guarantee of this procedure is that if $G$ has a minimum cut that is non-trivial, i.e., not composed of a single vertex, then $\lambda_{G}=\lambda_{G^{\prime}}$ with constant probability. The following theorem states this formally.
@@
The following technical lemma provides the guarantees of the $\tau$-star contraction procedure in a fully dynamic setting. Using this lemma our algorithms maintain a KT-style contraction of the graph $G$ as a weighted graph (by merging parallel edges) in a fully dynamic setting. We call a contraction $G^{\prime}$ of $G$ complete, if it maps every edge of $G$ to an edge in $G^{\prime}$, and otherwise it is incomplete. [^3] Notice that a incomplete contraction is not a contraction in the strict sense, since it does not map every edge of $G$ to an edge in $G^{\prime}$. Our contraction procedure only tracks edges that are internal to the center vertices $R$, hence if a vertex $v\in V\setminus R$ is not contracted, then its incident edges are not mapped in $G^{\prime}$. Therefore, its contraction is complete only when all vertices in $V\setminus R$ are contracted. Fortunately, this happens with high probability whenever $\tau\geq\delta_{G}$ by Theorem 2.2.
@@ 3.1 Amortized Time Complexity
$\Pr\left[R_{v}\text{ changes}\right]=\Pr\left[u\in R\right]\cdot\Pr\left[R_{v}\text{ changes}\mid u\in R\right]=p\cdot\Pr\left[R_{v}\text{ changes}\mid u\in R\right],$ (1)
@@ 3.2 From Amortized to Worst-Case Time Complexity via Incremental Rebuilding
$\Pr\left[\sum_{i=1}^{N}I_{i}\geq 1\right]\leq\frac{\mathbb{E}\left[\sum_{i=1}^{N}I_{i}\right]}{1}\leq O\left(\frac{1}{\log^{2}n}\right).\qed$

### v2

@@ 1.1 Contributions
Our second algorithm improves on the result of [GHN+23] for large edge connectivity, namely when $\lambda_{G}=\tilde{\omega}(\sqrt{n})$. Our result is actually stronger, since it is parametrized by the minimum degree of the graph at the time of the update, $\delta_{G}$, which is only larger than the edge connectivity $\lambda_{G}$ and hence it implies an update time of $\tilde{O}(n^{2}/\lambda^{2}_{G})$. Therefore, Theorem 1.2 improves on [GHN+23] for large edge connectivity, i.e. $\lambda_{G}\geq\tilde{\omega}(\sqrt{n})$, and is in fact the first algorithm to achieve worst-case update time $o(n)$ for large edge connectivity. Moreover, the improvement applies also to graphs with large minimum degree but small edge connectivity. The improvement is even larger when separating the update and query times, i.e. if the algorithm only needs to report the edge connectivity occasionally.
@@ 1.2 Future Work
The main open question remaining is to achieve worst-case update time $o(n)$ for all edge connectivity values. Prior work achieves this for edge connectivity $\lambda_{G}=o(n^{1/11}/\operatorname{polylog}n)$ [Tho07, JST24, dVC25]. Our results complement these results and achieve it for large edge connectivity, namely $\lambda_{G}=\tilde{\omega}(\sqrt{n})$. Perhaps surprisingly, the remaining case is that of intermediate connectivity, $\lambda_{G}\in[n^{1/11},n^{1/2}]$.
@@ 2 Proof of Main Theorems
Our algorithms follow the general recipe of [GHN+23], with the main difference being that our contraction algorithm maintains an explicit KT-style contraction as a weighted graph. [^2] To recover the edges of a minimum cut we can maintain a list of the edges of $G$ that are mapped to each edge of $G^{\prime}$ and report the edges corresponding to the minimum cut in $G^{\prime}$. This both simplifies the algorithm and improves the update time whenever $\delta_{G}\geq\omega(\sqrt{n})$. Specifically, we leverage the $\tau$-star contraction procedure of [AEG+22, KK25], which was originally designed for finding a global minimum cut in the cut-query model. The procedure works as follows. Sample a set of center vertices $R\subseteq V$ by including each vertex $v\in V$ independently with probability $p=O(\log n/\tau)$. Let $H={\{v\in V\setminus R\mid d_{G}(v)\geq\tau\}}$ be the set of non-center vertices with degree at least $\tau$, where throughout $d_{G}(v)$ is the degree of $v$ in $G$. Then, for every $v\in H$, uniformly sample a vertex $r\in N_{G}(v)\cap R$, where $N_{G}(v)$ is the neighborhood of $v$ in $G$, and contract the corresponding edge $(v,r)$, keeping parallel edges, which yields a contracted multigraph $G^{\prime}$. The main guarantee of this procedure is that if $G$ has a minimum cut that is non-trivial, i.e., not composed of a single vertex, then $\lambda_{G}=\lambda_{G^{\prime}}$ with constant probability. An illustration of the $\tau$-star contraction is given in Figure 2. The following theorem states this formally.
@@
Figure 2: An illustration of the $\tau$-star contraction procedure with $\tau=2$. (i) A random set of center vertices $R$ is sampled (indicated by red). (ii) Each non-center vertex in $H=\{v\in V\setminus R\mid d_{G}(v)\geq\tau\}$ chooses a random neighbor in $R$ (indicated by dashed edges); vertices not in $H\cup R$ are indicated by a striped pattern. (iii) The corresponding edges are contracted (keeping parallel edges) to obtain a contracted graph $G^{\prime}$.
The following technical lemma provides the guarantees of the $\tau$-star contraction procedure in a fully dynamic setting. Using this lemma our algorithms maintain a KT-style contraction of the graph $G$ as a weighted graph (by merging parallel edges) in a fully dynamic setting. We call a contraction $G^{\prime}$ of $G$ complete, if it maps every edge of $G$ to an edge in $G^{\prime}$, and otherwise it is incomplete. [^3] Notice that a incomplete contraction is not a contraction in the strict sense, since it does not map every edge of $G$ to an edge in $G^{\prime}$. Our contraction procedure only tracks edges that are internal to the center vertices $R$, hence if a vertex $v\in V\setminus R$ is not contracted, then its incident edges are not mapped in $G^{\prime}$. Therefore, its contraction is complete only when all vertices in $V\setminus R$ are contracted. Fortunately, this happens with high probability whenever $\delta_{G}\geq\tau$ by Theorem 2.2.
@@ 3.1 Amortized Time Complexity
(1) $\Pr\left[R_{v}\text{ changes}\right]=\Pr\left[u\in R\right]\cdot\Pr\left[R_{v}\text{ changes}\mid u\in R\right]=p\cdot\Pr\left[R_{v}\text{ changes}\mid u\in R\right],$
@@ 3.2 From Amortized to Worst-Case Time Complexity via Incremental Rebuilding
$\Pr\left[\sum_{i=1}^{N}I_{i}\geq 1\right]\leq\frac{\mathbb{E}\left[\sum_{i=1}^{N}I_{i}\right]}{1}\leq O\left(\frac{1}{\log^{2}n}\right).$